a small piece of wire bent into L shape with upright and horizontal positiions of equal lengths, is placed with horizontal position along axis of concave mirrror whose radius of curvature is 10cm. If bend is 20cm from pole of mirror, then ratio of lengths of images of upright and horizontal positions of wire is___________________.


 


ans: 3:1

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

1/u +1/v =1/f=2/r,f=5cm, 1/20 +1/v = 1/f5, v=20/3 cm, magnification=-v/u=-1/3,ratio=1/3